A company's costs, in millions of dollars, are given by the equation, C=x^2-3x-27, where x is the number of items sold, in thousands. What are the costs when 1,000 items are sold? and 1,500 items are sold?
So far I have used 1 and 1.5 and substituted into the equation but I am getting negative numbers is that correct?

Answer by stanbon(75887)

A company's costs, in millions of dollars, are given by the equation, C=x^2-3x-27, where x is the number of items sold, in thousands. What are the costs when 1,000 items are sold? and 1,500 items are sold?
-------------------
C(x) = x^2-3x-27
C(1) = 1-3-27 = -29
C(1.5) = 2.25 - 3*1.5 - 27 = 2.25 - 4.5 - 27 = -29.25
-----
Note: If x is in thousands, these answers are what you get.
As costs the answers do not make much sense. Maybe x
should be 1000 instead of 1 and 1500 instead of 1.5
That way your cost figures would be positive. Try it.
------------------
Cheers,
Stan H.